OpenStudy (anonymous):

Which of the following is a solution to the following equation y=3x-1

OpenStudy (anonymous):

The answers are (4,1) (2,5) (4,3) (0,3)

Directrix (directrix):

Take the equation and substitute in for the x and y values and see which one results in a true statement. For the first option, (4,1), x = 4 and y = 1. Do those values make y=3x-1 true? Does 1 = 3*4 - 1? Does 1 = 12 - 1 = 11? Does 1 = 11? No. Therefore, the answer is NOT (4,1).

Directrix (directrix):

@Briley_Dunbar123 You try the second one. (2,5) y=3x-1 Is 5 = 3*2 - 1 ?